Examine the function ,f(x)=(x+1)^3(x-3)^2 for extreme values

1
Expert's answer
2022-01-11T00:02:57-0500
"f(x)=(x+1)^3(x-3)^2"

Find the first derivative with respect to "x"


"f'(x)=3(x+1)^2(x-3)^2+2(x+1)^3(x-3)"

"=(x+1)^2(x-3)(3x-9+2x+2)"

"=(x+1)^2(x-3)(5x-7)"

Find the critical number(s)


"f'(x)=0=>(x+1)^2(x-3)(5x-7)=0"

Critical numbers: "-1, 1.4, 3."

If "x<-1,f'(x)>0, f(x)" increases.

If "-1<x<1.4,f'(x)>0, f(x)" increases.

If "1.4<x<3,f'(x)<0, f(x)" decreases.

If "x>3,f'(x)>0, f(x)" increases.


"f(-1)=(-1+1)^3(-1-3)^2=0"


"f(1.4)=(1.4+1)^3(1.4-3)^2=35.38944"

"f(3)=(3+1)^3(3-3)^2=0"

The function "f" has a local maximum with value of "35.38944" at "x=1.4."

The function "f" has a local minimum with value of at "x=3."

The function "f" has neither a local maximum nor a local minimum at "x=-1."